OpenStudy (ladylauren):

WILL FAN OR MEDAL!!!!!!!! The table and the graph below each show a different relationship between the same two variables, x and y:

OpenStudy (ladylauren):

How much more would the value of y be on the graph than its value in the table when x = 12?

OpenStudy (ladylauren):

20 30 60 70

OpenStudy (angelica):

In the first pic, 4 * n = 100 which means n = 25 hence 4 = 100 Find y using 12*25 and look at the graph how much difference does the answer is from the last point

OpenStudy (ladylauren):

so 300

OpenStudy (angelica):

n is equal to 25 because I divided 100 by 4

OpenStudy (angelica):

Yes, I believe so

OpenStudy (ladylauren):

so 20

OpenStudy (angelica):

To find the value you are looking for, you have to look how much distance it is in the middle of 300 and 240 as shown in the graph

OpenStudy (ladylauren):

that isn't a option

OpenStudy (flying_dutchman):

60! in the table x equals 25y, so 12x is equal to 300y on the graph x equals 30y, so 12x is equal to 360y then you only have to subtract 360-300

OpenStudy (ladylauren):

ok thx

OpenStudy (angelica):

It is. Look for the difference of 300 and 240 and that is the answer. How did you get 20?

OpenStudy (aaronandyson):

For an increase of 1 in the x table there is an increase of 25 in the y table.

OpenStudy (aaronandyson):

12*25 = 300

OpenStudy (aaronandyson):

To get the difference subtract 240 from 300

OpenStudy (aaronandyson):

300-240 = 60

OpenStudy (ladylauren):

k thx so much